Tap A takes 4 hours to fill a tank. Tap B takes 6 hours to fill the same tank. How long does the tank take to fill half if both tap was on at se te?
Answer by 303795(602)

Tap A fills the tank in 4 hours so it fills 1/4 of the tank each hour.
Tap B fills the tank in 6 hours so it fills 1/6 of a tank in one hour.
In one hour the first Tap A fills 1/4 of the tank and tap B fills another 1/6 of the tank i.e.
1/4 + 1/6 = 3/12 + 2/12 = 5/12 filled per hour.
This means that it would take 12/5 hours (or 2.4 hours)to fill the tank. It would be half filled in 1.2 hours ( or 1 hour and 12 (0.2 x 60)minutes)